for a triangle when c = 60 degree prove 1/b+c + 1/a+c = 3/a+b+c
Answer by cleomenius(959)

I assume you are working with an equilateral triangle.
b = 60 degrees. Given
a = 60 degrees The angles in an equilateral tringle are equal.
c = 60 degrees The angles in an equilateral triangle are equal.
1/b + c = 1/120 addition
1/a + c = 1/120 addition
1/120 + 1/120 = 2/120 = 1/60 addition, reduction.
3/a + b + c = 3/180 substitution
3/180 = 1/60 Reduction.
They are the same, 1/60.
Cleomenius.